A company produces three products which are interdependent. These are A, B and C. The flow of inputs and outputs between the products is represented in the table below:


Inputs (in thousands of units)


A B C Final demand


Outputs


(in thousands of units) A


B


C 40


60


80 65


130


65 75


75


25 20


60


80



Required:


i) Derive the technical coefficients matrix (3 marks)


ii) Determine the Leontief’s inverse matrix (12 marks)


iii) Compute the output level for each product if the final demand for product A increased by 2000 units, that of product C decreased by 1,000 units and the final demand for product B remained unchanged. (5 marks)





1
Expert's answer
2022-03-04T11:56:11-0500


i) The technical coefficients matrix

The technical coefficient matrix refers to the matrix of B.

B= "\begin{bmatrix}
b11 & b12 & b13 \\
b21 & b22 & b23 \\
b31 & b32 & b33 \\
\end{bmatrix}"

where,

b11 = 40/200 = 0.2

b12= 65/325= 0.2

b13= 75/250 = 0.3


b21=60/200 = 0.3

b22=130/325 = 0.4

b23=75/250 = 0.3


b31=80/200= 0.4

b32=65/325 = 0.2

b33= 25/250 = 0.1

hence,


B= "\begin{bmatrix}
0.2 & 0.2 & 0.3 \\
0.3 & 0.4 & 0.3 \\
0.4 & 0.2 & 0.1 \\
\end{bmatrix}" [Answer]


ii) The Leontief’s inverse matrix

Leontief's inverse matrix is given by (I-B)-1

I-B = "\begin{bmatrix}
1 &0 & 0 \\
0 & 1 & 0 \\
0 & 0 &1 \\
\end{bmatrix} - \begin{bmatrix}
0.2 & 0.2 & 0.3 \\
0.3 & 0.4 & 0.3 \\
0.4 & 0.2 & 0.1 \\
\end{bmatrix}=\begin{bmatrix}
0.8 & -0.2 & -0.3 \\
-0.3 & 0.6 & -0.3 \\
-0.4 & -0.2 & 0.9 \\
\end{bmatrix}"


(I-B)-1

"\begin{bmatrix}
0.8 & -0.2 & -0.3 \\
-0.3 & 0.6 & -0.3 \\
-0.4 & -0.2 & 0.9 \\
\end{bmatrix}|\begin{bmatrix}
1 &0 & 0 \\
0 & 1 & 0 \\
0 & 0 &1 \\
\end{bmatrix}" Turn the left matrix into the right matrix


R1  ( "\frac{5}{4}" )R1


"\begin{bmatrix}
1 & \frac{-1}{4} & \frac{-3}{8} \\
\frac{-3}{10} & \frac{3}{5} & \frac{-3}{10} \\
\frac{-2}{5} &\frac{-1}{5} & \frac{9}{10} \\
\end{bmatrix}|\begin{bmatrix}
\frac{5}{4} &0 & 0 \\
0 & 1 & 0 \\
0 & 0 &1 \\
\end{bmatrix}"


R2 R2- ( "\frac{-3}{10}" )R1


"\begin{bmatrix}
1 & \frac{-1}{4} & \frac{-3}{8} \\
0 & \frac{21}{40} & \frac{-33}{80} \\
\frac{-2}{5} &\frac{-1}{5} & \frac{9}{10} \\
\end{bmatrix}|\begin{bmatrix}
\frac{5}{4} &0 & 0 \\
\frac{3}{8} & 1 & 0 \\
0 & 0 &1 \\
\end{bmatrix}"


R3  R3 -"(\\displaystyle - \\frac{2}{5}\n\n\u2212\u200b5\n\n\u200b\n\n\u200b2\n\n\n\n\u200b\u200b)R1"


"\begin{bmatrix}
1 & \frac{-1}{4} & \frac{-3}{8} \\
0 & \frac{21}{40} & \frac{-33}{80} \\
0 &\frac{-3}{10} & \frac{3}{4} \\
\end{bmatrix}|\begin{bmatrix}
\frac{5}{4} &0 & 0 \\
\frac{3}{8} & 1 & 0 \\
\frac{1}{2} & 0 &1 \\
\end{bmatrix}"


R2  "(\displaystyle 1.90476)R2"


"\begin{bmatrix}
1 & \frac{-1}{4} & \frac{-3}{8} \\
0 & 1 & \frac{-11}{14} \\
0 &\frac{-3}{10} & \frac{3}{4} \\
\end{bmatrix}|\begin{bmatrix}
\frac{5}{4} &0 & 0 \\
\frac{3}{8} & \frac{40}{21} & 0 \\
\frac{1}{2} & 0 &1 \\
\end{bmatrix}"


R3  R3 - "(\\displaystyle - \\frac{3}{10}\u200b\u200b)R2"


"\begin{bmatrix}
1 & \frac{-1}{4} & \frac{-3}{8} \\
0 & 1 & \frac{-11}{14} \\
0 &0 & \frac{18}{35} \\
\end{bmatrix}|\begin{bmatrix}
\frac{5}{4} &0 & 0 \\
\frac{3}{8} & \frac{40}{21} & 0 \\
\frac{5}{7} & \frac{4}{7} &1 \\
\end{bmatrix}"



R3  "(\displaystyle 1.94444)R3"


"\begin{bmatrix}
1 & \frac{-1}{4} & \frac{-3}{8} \\
0 & 1 & \frac{-11}{14} \\
0 &0 & 1 \\
\end{bmatrix}|\begin{bmatrix}
\frac{5}{4} &0 & 0 \\
\frac{3}{8} & \frac{40}{21} & 0 \\
\frac{25}{18} & \frac{10}{9} &\frac{35}{18} \\
\end{bmatrix}"


R2  R2 - "(\displaystyle -0.785714)R3"


"\begin{bmatrix}
1 & \frac{-1}{4} & \frac{-3}{8} \\
0 & 1 & 0\\
0 &0 & 1 \\
\end{bmatrix}|\begin{bmatrix}
\frac{5}{4} &0 & 0 \\
\frac{65}{36} & \frac{25}{9} & \frac{55}{36} \\
\frac{25}{18} & \frac{10}{9} &\frac{35}{18} \\
\end{bmatrix}"


R1  R1 - "(\\displaystyle - \\frac{3}{8}\u200b\u200b)R3"


"\begin{bmatrix}
1 & \frac{-1}{4} & 0 \\
0 & 1 & 0\\
0 &0 & 1 \\
\end{bmatrix}|\begin{bmatrix}
\frac{85}{48} &\frac{5}{12} & \frac{35}{48} \\
\frac{65}{36} & \frac{25}{9} & \frac{55}{36} \\
\frac{25}{18} & \frac{10}{9} &\frac{35}{18} \\
\end{bmatrix}"


R1  R1 - (\displaystyle - \frac{1}{4}​​)R2


"\begin{bmatrix}
1 & 0 & 0 \\
0 & 1 & 0\\
0 &0 & 1 \\
\end{bmatrix}|\begin{bmatrix}
\frac{20}{9} &\frac{10}{9} & \frac{10}{9} \\
\frac{65}{36} & \frac{25}{9} & \frac{55}{36} \\
\frac{25}{18} & \frac{10}{9} &\frac{35}{18} \\
\end{bmatrix}"


Therefore,

[I-B]-1= "\begin{bmatrix}
\frac{20}{9} &\frac{10}{9} & \frac{10}{9} \\
\frac{65}{36} & \frac{25}{9} & \frac{55}{36} \\
\frac{25}{18} & \frac{10}{9} &\frac{35}{18} \\
\end{bmatrix}" [Answer]


iii) Compute the output level for each product if the final demand for product A increased by 2000 units, that of product C decreased by 1,000 units and the final demand for product B remained unchanged.

X= [I-B]-1D


D="\begin{bmatrix}
20+2 \\
60\\
80-1\\
\end{bmatrix} =
\begin{bmatrix}
22 \\
60\\
79\\
\end{bmatrix}"


X="\begin{bmatrix}
\frac{20}{9} &\frac{10}{9} & \frac{10}{9} \\
\frac{65}{36} & \frac{25}{9} & \frac{55}{36} \\
\frac{25}{18} & \frac{10}{9} &\frac{35}{18} \\
\end{bmatrix} \begin{bmatrix}
22 \\
60\\
89\\
\end{bmatrix}"


Multiply the rows of Matrix A with the columns of Martrix B


X= "\begin{bmatrix}
\frac{20.22}{9} +\frac{10.60}{9} + \frac{10.79}{9} \\
\frac{22.65}{36} + \frac{55.79}{36} + \frac{25.60}{9} \\
\frac{25.25}{18} + \frac{10.60}{9} +\frac{35.79}{18} \\
\end{bmatrix}"


="\begin{bmatrix}
\frac{610}{3} \\
\frac{3925}{12} \\
\frac{1505}{6} \\
\end{bmatrix}" (In thousands of units) [Answer]